OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. The value of g(-2) is smaller than the value of g(4). The value of g(-2) is the same as the value of g(4). The values of g(-2) and g(4) cannot be compared. The value of g(-2) is larger than the value of g(4).

OpenStudy (anonymous):

I think it is A

OpenStudy (anonymous):

Am I Correct?

OpenStudy (anonymous):

I just plugged in the -2 and 4 into the g...

jhonyy9 (jhonyy9):

g(-2)=(-3)(-2)+4= 10 yes ?

OpenStudy (anonymous):

what

jhonyy9 (jhonyy9):

g(4)=-3*4 +4 = -12+4=8

jhonyy9 (jhonyy9):

so then how you think ?

jhonyy9 (jhonyy9):

the last one

jhonyy9 (jhonyy9):

yes ?

OpenStudy (anonymous):

OK... Thank you..